Cristiano Ronaldo shows off his skill with a soccer ball by kicking it 45.20 m
m
into a bucket. He kicks the ball at 37.00 ∘

above horizontal, and the bucket is placed at the same level from which the ball is kickeWhat initial speed does he need to kick the ball at so it lands in the bucket?
1
Expert's answer
2020-06-08T10:25:10-0400

The range


"R=\frac{v_0^2\sin2\theta}{g}"

Hence, the initial speed of a ball


"v_0=\sqrt{\frac{gR}{\sin2\theta}}\\
=\sqrt{\frac{9.81\times 45.20}{\sin(2\times 37.00^{\circ})}}=21.48\:\rm m/s"